Status Report

ISS On-Orbit Status 24 May 2002

By SpaceRef Editor
May 24, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below. Day 169 on orbit for Expedition 4.

After wake-up at 2:00 am EDT and the standard (mandatory non-work)
90-min. post-sleep and breakfast period, the crew held the regular
daily planning conference (via S-band) and prepared for today’s

CDR Yuri Onufrienko and FE-1 Carl Walz performed troubleshooting with
removal and replacement of the Potok-150MK (150 micron) air filter
unit of the SOGS air revitalization subsystem in the SM.

FE-2 Dan Bursch meanwhile had another session with the BPS (biomass
production system) test facility, performing gas sampling on PGC3
(plant growth chamber #3).  After he had closed the BPS carbon
dioxide (CO2) cylinder valve on 5/22, the residual CO2 in the lines
was allowed to drain off for 48 hours. Today, right after the gas
sampling, he changed out the CO2 cylinder for the BPS, after which
the ground did some commanding.  Later, he harvested wheat
plants in PGC1, the second harvest for this chamber.  A total of
four plants were collected for fixation, and two of them were stowed
in the ARCTIC freezer/refrigerator.  In order to minimize
condensation in the freezer chamber, Dan placed the fixed samples
inside a bag prior to stowage.

The two flight engineers completed the scheduled EVA plan review and
discussed it with the ground on S-band.  Preparations are
underway for the new pre-breathing procedure, which uses Shuttle O2
while doing the prescribed exercise on the cycle ergometer with a
breathing mask attached via a 90-ft hose to a Shuttle O2 port (LEH);
and for a modified power connection between the MT (mobile
transporter) and MBS (mobile base system).  [The subsequent
non-exercise portion of the mandatory 80-min. mask prebreathe is done
using O2 from the Airlock (A/L) PBA (portable breathing apparatus)
outlet via a 60-ft hose.  The EMU spacesuits themselves still
will get all their O2 for purge, EMU prebreathe and post-EVA recharge
from the A/L HPGTs (high-pressure gas tanks), but the selected
modified mask prebreathe procedure will yield an expected 80%
reduction of HPGT O2 consumption.  The second major EVA prep
task concerns the building of two new power crossover jumper cables
between MT and MBS, after it was discovered the originally planned
method of primary and backup connection would create undesirable
loops between the electrical return paths of the two redundant power

Onufrienko spent some more time with return cargo preparations
and pre-packing for UF-2.

All crewmembers performed their daily physical exercise program on
TVIS (treadmill) and RED (resistive exercise device).  Since the
pulley cables of the RED are past their cycle certification and there
is not enough time left before UF-2 to replace them, the ground
requested that crewmembers perform a canister cable inspection during
the first five minutes of daily RED exercise.  They were also
asked to wear safety straps during exercise.  The cable swap-out
will be done after UF-2.

FE-2 Dan Bursch completed today’s planned downloading of the exercise
files to the MEC (medical equipment computer), a periodic 40-min.

Walz reconfigured the CCAA (common cabin air assembly) in the Lab,
switching it from the starboard channel to the alternate port
channel.  [The CCAA is a network of ducting that draws in the
air through filters, delivers it for conditioning, and returns it to
the modules.]

He also power-cycled the IV-CPDS (intravehicular charged particle
directional spectrometer) off/on to set it up for ground commanding,
nominally done on Wednesdays and Fridays.

The regular daily routine maintenance was performed by Onufrienko
(SOSH life support systems servicing, BRPK-2 water separator
inspection, and IMS delta file preparation), and Carl Walz (Inc. 4/8A
payloads status checks.

Carl also completed his weekly FFQ (food frequency questionnaire)
form, which determines the nutritional value of the crew’s diet each
week, required to assess overall crew health.  The FFQ is
designed to provide an estimate of dietary intake without requiring
detailed logs of every item eaten.

A conjunction with a Russian SL-11 rocket body (object #3019) is
predicted for 5/26 (Sunday), with TCA (time of closest approach) at
3:48 pm EDT.  Current estimated miss distance: 9.7 km. The
object is being monitored very closely, and if a maneuver should be
necessary, preliminary TIG (time of ignition) will be on Sunday at
1:30 pm EDT.

Today’s CEO (crew earth observation) targets were Angolan Biomass
(as the broad bulge of the coast [Angola-Namibia
border] came into view, the crew was to look left and document any
smoke moving offshore from Angolan agricultural lands. Normally
exposed views against the background of the sea will permit estimates
of smoke thickness; including the coastline in the frame will aid in
locating/mapping any plumes)
, Mouth of Amazon (Dynamic
Event: The rainy season is getting under way in the Amazon drainage
basin; the great river and its tributaries will be carrying
progressively greater sediment loads. This pass allowed for a
companion view to one that was suggested earlier this week. The
Amazon delta and river mouth was to the station’s left, and they were
to record water color and plumes of sediment being delivered to the
Atlantic. Beyond the subaerial delta was a broad apron of mud, and
they were asked to photograph the extent of the muddy mass)
Parana River (there was a break in the cloud cover at
present, and the lowlands of the central Parana basin should have
been fully visible: the Pantanal area was left of track and the Ibera
swamp was right. Harvest should be complete so croplands are brown in
contrast with the natural cover. Paranapanema reservoir was on the
station’s left, in the Parana headwaters in the coast ranges; crew
was to watch for any evidence of dam construction and/or new
reservoirs NW of Florianopolis).

U.S. and Russian Segment Status (as of 1:52 pm EDT):

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode).
Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump
failed).  U.S. CDRA CO2 scrubber is Off.  BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify

SM Working Compartment:  Pressure (mmHg) — 757,
temperature (deg C) — 28.5, ppO2 (mmHg) — 146.9, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 762,
temperature (deg C) — 21.2.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 23.3.
Node:  Pressure (mmHg) — 748.98, temperature (deg C) —
24.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 751.18, temperature (deg C)
— 24.7, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 750.98,
temperature (deg C) — 24.8; shell heater temp (deg C) — 24.9, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 23.1
PMA-2:  Shell heater temp (deg C) — 11.7

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Autotrack mode.
SM batteries: Battery #2 is offline (ROM mode). All other batteries
(7) are in “partial charge” mode.
FGB: Battery #3 is offline (ROM mode).  All other batteries (5)
are in “partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating

SSRMS/Canadarm2 at Progress viewing position, with Keep Alive
power on both strings.
RWS (robotics workstations) are Off.
SSRMS Prime string Wrist Roll (WR) joint to be replaced on UF-2.

ISS Orbit (as of this morning, 6:47 am EDT [=
Mean altitude — 388.2 km
Apogee — 392.7 km
Perigee — 383.6 km
Period — 92.3 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0006691
Orbits per 24-hr. day — 15.60
Altitude decrease — 170 m (mean) in last 24 hours
Solar Beta Angle — 64.8 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 20017
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: ~180 deg, pitch: -5.6 deg., roll: 0

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.